OpenStudy (anonymous):

When asked to find d2y/dx2 implicitly in terms of x and y for x^2-y^2=9, I keep getting (y^2-x^2)/y^2 for my y'' and it should be (x^2-y^2)/y^2. What am I doing wrong?!

hartnn (hartnn):

lets check, \(\Large 2x-2yy' =0 \\ \large y' = \dfrac{x}{y}\)

hartnn (hartnn):

\(\Large y'' = \dfrac{y-xy'}{y^2}\) but \(y' =\dfrac{x}{y} \) so, \(\Large y'' = \dfrac{y-x(x/y)}{y^2} = \dfrac{y^2-x^2}{y^3}\) why do u sat the answer should be (x^2-y^2)/y^2. ??

hartnn (hartnn):

*say

hartnn (hartnn):

alternate way, \(\Large x=yy' \\ \large 1 = y'^2 +yy'' \\ \Large 1= x^2/y^2 +yy'' \\ \Large 1-x^2/y^2 = yy'' \\ \large y'' = \dfrac{y^2-x^2}{y^3}\) getting the same answer again!, so it must be correct :P

OpenStudy (anonymous):

because the original function was x^2-y^2 in the numerator. For the problem to work implicitly so that I can substitute 9 back into my y'', the numerator must equal x^2-y^2

hartnn (hartnn):

why *must* ?? if the numerator is y^2-x^2, we can always say that since, x^2-y^2 = 9 y^2-x^2 = -9 and our y" = -9/y^3

OpenStudy (anonymous):

Okay, I just wasn't sure if I could do that, but it is what I wanted to do all along, thanks

hartnn (hartnn):

all doubts cleared ? anything left anywhere ?

OpenStudy (anonymous):

No, we just never manipulated the original function when substituting before. I wasn't sure if it was safe to do. I'm good now.

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Thanks again